# DAL: A Practical Prior-Free Black-Box Framework for Non-Stationary Bandit Environments

### v2

@@ Abstract
We introduce a practical, black-box framework termed Detection Augmenting Learning (DAL) for the problem of non-stationary bandits without prior knowledge of the underlying non-stationarity. DAL is modular, accepting any stationary bandit algorithm as input and augmenting it with a change detector. Our approach is applicable to all common parametric and non-parametric bandit variants. Extensive experimentation demonstrates that DAL consistently surpasses current state-of-the-art methods across diverse non-stationary scenarios, including synthetic benchmarks and real-world datasets, underscoring its versatility and scalability. We provide theoretical insights into DAL’s strong empirical performance on piecewise stationary and drift settings, complemented by thorough experimental validation.
@@ 1 Introduction
Bandit models are widely applied in engineering domains, including recommendation systems [35, 36], online advertising [11, 48, 49], dynamic pricing [54], and real-time bidding [23]. Introduced through the multi-armed bandit (MAB) problem [42], many variants have since emerged, which fall into parametric bandits (PB) [4, 19, 22] and non-parametric bandits (NPB) [14, 52]. In a bandit problem, an agent selects actions $A_{t}$ from a compact set $\mathcal{A}\subseteq\mathbb{R}^{d}$ over $T$ rounds, receiving rewards
$X_{t}=f_{t}(A_{t})+\varepsilon_{t},$
@@
where $f_{t}:\mathcal{A}\rightarrow\mathbb{R}$ is the reward function and $\varepsilon_{t}$ is zero-mean sub-Gaussian noise. The goal is to minimize the dynamic regret, using a causal policy $\pi_{t}$ based on past interactions:
$R_{T}:=\mathbb{E}_{A_{t}\sim\pi_{t}}\left[\sum_{t=1}^{T}\max_{a\in\mathcal{A}}f_{t}(a)-f_{t}(A_{t})\right].$
In the case of PB, $f_{t}(A_{t})=\mu\left(\langle\theta_{t},A_{t}\rangle\right)$ where $\theta_{t}$ is an underlying (bounded) parameter unknown to the agent, and $\mu:\mathbb{R}\to\mathbb{R}$ is an injective function. This model includes linear bandits (LBs) with $\mu$ as identity, generalized linear bandits (GLBs), and self-concordant bandits (SCBs) where $\mu$ has the self-concordant property and the noise variance depends on the mean reward [44].
Regarding NPB, we consider the kernelized bandit (KB), under which $f_{t}$ belongs to a reproducing kernel Hilbert space (RKHS) $H_{k}$ induced by a continuous positive semi-definite kernel $k:\ \mathcal{A}\times\mathcal{A}\rightarrow\mathbb{R}$ with $k(x,x)\leq 1$ for all $x\in\mathcal{A}$ and $\|f_{t}\|_{H_{k}}\leq B$. An important quantity for KBs is the maximum information gain (MIG). The MIG at time $t$ is defined as $\gamma_{t}:=\max_{D\subset\mathcal{A}:|D|=t}I(y_{D};f_{D})$, where $I(y_{D};f_{D})$ denotes the mutual information between $f_{D}=[f(x)]_{x\in D}$ and $y_{D}=f_{D}+\varepsilon$, with $\varepsilon\sim\mathcal{N}(0,\lambda v^{2}I)$. It quantifies the reduction in uncertainty about $f$ after observing $y_{D}$ at points $D\subset\mathcal{A}$. For a compact subset $\mathcal{A}\subset\mathbb{R}^{d}$, $\gamma_{T}$ is $\mathcal{O}((\log T)^{d+1})$ for the Squared Exponential (SE) kernel, and $\mathcal{O}(T^{\frac{d(d+1)}{2\nu+d(d+1)}}\log T)$ for the Matérn kernel (with smoothness parameter $\nu$) [53].
@@
The lion’s share of the literature on bandits assumes stationarity—i.e., fixed $f_{t}$ or $\theta_{t}$—but this rarely holds in practice due to evolving user behavior, market conditions, etc.[3, 8, 12, 39, 51]. Non-stationarity is often categorized into two types–gradual drifts and abrupt changes. In the drift model, $f_{t}$ evolves slowly under a variation budget constraint [6, 58]. In contrast, piecewise stationary (PS) models assume abrupt shifts at unknown change-points:
$1=:\nu_{0}<\nu_{1}<\dots<\nu_{N_{T}}<\nu_{N_{T}+1}:=T+1,$
with $f_{t}=f_{t^{\prime}}$ for $t,t^{\prime}\in[\nu_{k},\nu_{k+1})$ and $f_{\nu_{k}-1}\neq f_{\nu_{k}}$ across change-points. Here, $N_{T}$ denotes the total number of changes.
The primary objective of this work is to develop a practical algorithm for non-stationary (NS) bandits that operates without prior knowledge of the underlying non-stationarity. In many real-world applications, changes in the environment—such as shifts in user behavior, emerging trends, or competitive interventions—can occur abruptly rather than gradually. These sudden and decisive shifts can severely impact algorithmic performance, whereas slow, continuous drifts often result in less dramatic degradation and can be handled more gracefully by existing methods. This makes the PS setting not only a natural modeling choice for non-stationarity, but also a case of significant practical relevance. Indeed, as we demonstrate in Section 4.3, algorithms designed for the PS regime can often extend their effectiveness to broader drifting scenarios. For this reason, our focus is firmly placed on the PS setting.
@@
NS bandit algorithms are typically either adaptive—adjusting actions continuously—or restarting, choosing to unlearn and kickstart the learning process at certain times. They may also be prior-based (assuming knowledge of the non-stationarity) or prior-free. Discounted and sliding-window algorithms that weigh recent observations more heavily and forget the distant past are examples of prior-based adaptive approaches, e.g., see [24, 30] for NS-MABs, [13, 45, 57] for NS-LBs, [21, 43, 57] for NS-GLBs, [44, 57] for NS-SCBs, and [16, 63] for NS-KBs. The budget-restarting approaches in [6] for NS-MABs, [61] for NS-LBs, [61] for NS-GLBs, and [63] for NS-KBs are examples of prior-based restarting algorithms. Finally, detection-based restarting algorithms come in both prior-based and prior-free forms. For NS-MABs, [9, 38] constitute such prior-based methods, while [5, 7, 27] are examples of prior-free approaches. For NS-LBs and NS-KBs, see [26]. The most closely related work is that of [27], which addresses the PS-MAB setting and introduces techniques that we build upon in establishing our theoretical insights. However, our setting is much more complex, as it extends to possibly infinite action spaces, where the ideas of [27] are not straightforward to apply. Our framework also applies to the NS-MAB setting, since the NS-LB reduces to NS-MAB as a special case [33]. However, we do not focus on this setting due to the existence of prior approaches [7, 27].
Among prior-free methods, black-box approaches are particularly appealing: they equip any stationary bandit algorithm with non-stationarity handling capabilities. MASTER [58] is the only known order-optimal black-box method for general reinforcement learning. However, as shown in [25], MASTER’s detection mechanism fails to trigger in realistic settings, leading to poor practical performance in experiments. Our work addresses this limitation through the proposal of a black-box framework that is able to achieve good practical performance.
Our Contributions
We introduce a practical, prior-free, detection-based restarting algorithm that is tailored for non-stationary bandits—the first of its kind, to our knowledge. This design choice reflects three key practical insights: (i) prior knowledge of non-stationarity is rarely available; (ii) detection-based methods often enjoy lower worst-case complexity compared to adaptive approaches [40]; and (iii) the black-box paradigm results in significant simplification by reducing the design of algorithms for the non-stationary setting to that of designing how to restart stationary algorithms. In particular, we use a change detector that works with samples obtained from forced exploration to trigger restarts, and combine that with a stationary bandit algorithm. Our design is modular, easy to implement, offering flexibility and adaptability. Extensive empirical evaluation on both synthetic and real-world non-stationary benchmarks demonstrates that our approach consistently outperforms existing state-of-the-art approaches, both prior-free and prior-based. Finally, we provide theoretical insights on the performance of our framework. We show that in the PS setting, under mild conditions, our algorithm’s regret bound matches the state-of-art in PS-LBs and PS-GLBs and improves the current best known bounds for PS-SCBs and PS-KBs. Regarding drifting environments, we offer an intuitive explanation for the conditions under which our approach can achieve good performance, which is validated through additional experiments.
@@ 2 The DAL Framework
To detect changes in the environment, DAL monitors shifts in the representative statistics of the environment, using a general-purpose detector $\mathcal{D}$ for changes in sequences of independent random variables. This distinguishes our approach from methods like MASTER [58], which rely on detecting violations of stationary regret guarantees—an approach not suited for broader detection tasks. In bandit problems, the environment is characterized by reward distributions across actions, so effective change detection requires identifying shifts in these distributions–a premise that strongly correlates with good practical performance, according to [25]. The art lies in employing a detection mechanism that resets learning shortly after a change-point $\nu$, while suppressing the probability of false alarm. We adopt a detection method aligned with [7, 27], grounded in the well-established theory of quickest change detection [41, 55, 56, 59]. DAL keeps a sample history for the rewards of actions that belong in a carefully selected finite subset, and aims to detect changes in those actions.
@@
In stationary bandit settings, effective algorithms quickly concentrate on (near-)optimal actions, rarely exploring suboptimal ones. In NS environments, however, this behavior may lead to missed changes on rarely sampled suboptimal actions. To detect such changes, forced exploration is essential. When the action space is large or infinite, uniformly exploring all actions becomes infeasible. To handle infinite action spaces, DAL is able to abstract the detection task through a finite covering set, $\mathcal{A}_{\mathrm{e}}=\{a^{(i)}:i\in[N_{\mathrm{e}}]\}\subseteq\mathcal{A}$, where we specifically use $a^{(i)}$ to denote the $i$-th action in $\mathcal{A}_{\mathrm{e}}$. $\mathcal{A}_{\mathrm{e}}$ must be selected in a way that can guarantee that all significant changes in the reward function can be identified. To do this, the designed $\mathcal{A}_{\mathrm{e}}$ must have at least one action $a^{(i)}$ such that $f_{t}(a^{(i)})$ has changed. DAL is forced to play each action in $\mathcal{A}_{\mathrm{e}}$ once in the first $N_{\mathrm{e}}$ rounds, and then follows the stationary bandit algorithm for the next $\lceil N_{\mathrm{e}}/\alpha_{k}\rceil-N_{\mathrm{e}}$ steps. Here, $\alpha_{k}$ controls the exploration frequency, enabling a balance between performance and reliable change detection.
@@
The core idea of DAL is to augment any bandit algorithm $\mathcal{B}$ with a change detector $\mathcal{D}$. To facilitate accurate detection, DAL periodically forces $\mathcal{B}$ to sample from the predefined exploration (covering) set $\mathcal{A}_{\mathrm{e}}$. If $\mathcal{D}$ detects a change in the reward distribution associated with any action in $\mathcal{A}_{\mathrm{e}}$, DAL resets the bandit algorithm $\mathcal{B}$ entirely. A key advantage of DAL is its ability to translate strong stationary performance into robust performance under NS conditions. Therefore, by selecting a well-performing stationary algorithm, the DAL framework inherently achieves effective adaptation to NS environments. In fact, the only requirement for DAL’s input stationary algorithm is for it to attain optimal stationary regret performance bounds.
@@
Input: bandit $\mathcal{B}$, detector $\mathcal{D}$, covering set $\mathcal{A}_{\mathrm{e}}$, action set $\mathcal{A}$, horizon $T$, frequencies $\{\alpha_{k}\}_{k=1}^{\infty}$
Initialize: detection $\tau\leftarrow 0$, histories $\mathcal{H}_{a,\mathcal{D}}\leftarrow\emptyset$ $\forall\,a\in\mathcal{A}_{\mathrm{e}}$, counter $k\leftarrow 1$
@@
2:     if $(t-\tau-1)\mod\lceil N_{\mathrm{e}}/\alpha_{k}\rceil+1=i\leq N_{\mathrm{e}}$ then
3:        Play action $a^{(i)}$ and receive the reward $X_{t}$.
4:        Add sample $X_{t}$ into history $\mathcal{H}_{a^{(i)},\mathcal{D}}$.
5:        if $\mathcal{D}\left(\mathcal{H}_{a^{(i)},\mathcal{D}}\right)=\mathrm{detection}$ then
6:           Reset the bandit algorithm $\mathcal{B}$.
7:           Empty sample histories for all actions in $\mathcal{A}_{\mathrm{e}}$.
8:           $\tau\leftarrow t$, $k\leftarrow k+1$
9:        end if
10:     else
11:        Run the stationary bandit algorithm $\mathcal{B}$.
12:     end if
13:  end for
@@ 3 Practical Performance
In this section, our goal is to establish DAL’s practical effectiveness, by providing experiments on piecewise stationarity and drifting non-stationarity for both synthetic and real-world benchmarks. Since no existing method directly targets the PS-PB and PS-NPB settings with order-optimal guarantees, we compare DAL against the most relevant state-of-the-art algorithms adaptable to PS and drifting settings. We first include MASTER [58], the only other black-box algorithm with order-optimal regret bounds. Although MASTER lacks formal guarantees for SCBs, its recent empirical evaluations [57] support combining it with Logistic-UCB-1 [19] for NS-SCBs. For NS-KBs, despite theoretical results indicating MASTER could be combined with GPUCB [14], MASTER does not attain the state-of-the-art regret bound [26], and due to existing practical issues [25], we omit it from our experiments in NS-KBs. We also evaluate ADA-OPKB [26], the state-of-the-art, prior-free, order-optimal method for NS-LBs and NS-KBs. ADA-OPKB requires extensive parameter tuning for strong performance, posing a challenge in a truly prior-free environment. For fair comparison, however, we tune its parameters to optimize its performance in our evaluation. Finally, we benchmark DAL against two leading prior-based discounted methods: WeightUCB [57], with best theoretical and practical performance for drifting PBs and PS-SCBs, and WGP-UCB [17] for its good performance in drifting KBs. All algorithms use recommended parameters from their original works.
Regarding DAL’s input, in all settings we employ the Generalized Likelihood Ratio (GLR) test [7, 27, 29] as the detector $\mathcal{D}$. Specifically, for the NS-LB setting, we combine DAL with LinUCB [1] and the Gaussian version of the GLR test. For the NS-GLB setting, we employ DAL with GLM-UCB [22] and the Gaussian GLR test. For the NS-SCB setting, we employ DAL with OFUGLB [34], along with the Bernoulli GLR test. Finally, for the NS-KB setting, we combine DAL with REDS [47] along with the Gaussian GLR test. To identify the cover set $\mathcal{A}_{\mathrm{e}}$, we use the information provided in Section 4.1. To implement and practically tune the detectors, we follow the information provided in [27], and provide their structure in the Appendix. Finally, for all settings, we set $\alpha_{k}=10^{-2}\sqrt{kN_{\mathrm{e}}\log(T)/T}$.

### v4

# DAL: A Practical Prior-Free Black-Box Framework for Non-Stationary Bandits

@@ Abstract
We introduce a practical, black-box framework termed Detection Augmented Learning (DAL) for the problem of non-stationary bandits without prior knowledge of the underlying non-stationarity. DAL accepts any stationary bandit algorithm as input and augments it with a change detector, enabling applicability to all common bandit variants. Extensive experimentation demonstrates that DAL consistently surpasses current state-of-the-art methods across diverse non-stationary scenarios, including synthetic benchmarks and real-world datasets, underscoring its versatility and scalability. We provide theoretical insights into DAL’s strong empirical performance, complemented by thorough experimental validation.
@@ 1 Introduction
Bandit models underpin a wide range of engineering systems, from recommendation and ads to dynamic pricing and real-time bidding [39, 40, 53, 54, 58, 25]. Many variants of multi-armed bandits (MABs) have emerged since the work of  [47], which fall into parametric bandits (PB) [6, 21, 24], non-parametric bandits (NPB) [57] and contextual bandits (CB) [63, 36]. In the general bandit problem, in each round, an agent receives a context $C_{t}$ randomly sampled from a set $\mathcal{C}$, and selects a policy $\pi_{t}$ from a policy set $\Pi$—a set of mappings from $\mathcal{C}$ to a compact action set $\mathcal{A}\subseteq\mathbb{R}^{d}$. Then, the agent chooses action $A_{t}=\pi_{t}(C_{t})$ and receives reward
$X_{t}=f_{t}(C_{t},A_{t})+\varepsilon_{t},$
@@
where $f_{t}:\mathcal{C}\times\mathcal{A}\rightarrow\mathbb{R}$ is the reward function and $\varepsilon_{t}$ is the zero-mean sub-Gaussian noise. The goal is to minimize the dynamic regret, using a causal policy $\pi_{t}$ based on past interactions:
$R_{T}:=\mathbb{E}_{\begin{subarray}{c}A_{t}\sim\pi_{t}\\ C_{t}\sim\mathcal{P}_{t}\end{subarray}}\left[\sum_{t=1}^{T}\max_{\pi\in\Pi}f_{t}(C_{t},\pi(C_{t}))-f_{t}(C_{t},A_{t})\right].$
[…]
@@
Bandits remain practically relevant today: recent deployments span A/B testing [66], clinical trials [59], large language models [56], diffusion models [5], and computer architecture [29], which even leverage the canonical formulations as the core decision engine. Accordingly, the key challenge is developing bandit methods that perform reliably under real-world constraints—aimed at practical effectiveness, not just analysis. The lion’s share of the literature on bandits assumes stationarity—i.e., fixed $f_{t}$, $\theta_{t}$, $\mathcal{P}_{t}$—but this rarely holds in practice due to evolving conditions [4, 11, 14, 44]. Non-stationary (NS) settings are often categorized into two types–gradual drifts and abrupt changes. In the drifting model, $f_{t}$ and $\mathcal{P}_{t}$ evolve slowly under a variation budget constraint [8, 62]. In contrast, piecewise stationary (PS) models assume abrupt shifts at unknown change-points:
$1=:\nu_{0}<\nu_{1}<\dots<\nu_{N_{T}}<\nu_{N_{T}+1}:=T+1,\quad N_{T}:\text{total number of changes}$
with $f_{t}=f_{t^{\prime}}$ and $\mathcal{P}_{t}=\mathcal{P}_{t^{\prime}}$ for $t,t^{\prime}\in\{\nu_{k},\dots,\nu_{k+1}-1\}$ and different across change-points.
@@
NS bandit algorithms are typically either adaptive—adjusting continuously, or restarting—choosing to unlearn and kickstart the learning process at certain times. They may also be prior-based (assuming knowledge of the non-stationarity) or prior-free. Prior-based adaptive methods (discounting/sliding window) weigh recent observations more heavily: NS-MABs [27, 33], NS-LBs [16, 50], NS-GLBs [23, 48], NS-SCBs [49, 61], NS-KBs [19, 70]. Prior-based restarting approaches use budgeted restarts: NS-MABs [8], NS-LBs/GLBs [68], NS-KBs [70]. Detection-based restarting methods exist in both flavors: prior-based for NS-MABs [12, 43] and NS-CBs [45]; prior-free for NS-MABs [7, 9, 31], for NS-LBs/KBs [30] and for NS-CBs [64, 15]. The most closely related work is that of [31], which addresses PS-MABs and introduces techniques that we build upon in establishing our theory. However, our setting is much more complex, as it extends to general NS bandits.
Among prior-free methods, black-box approaches are particularly appealing: they equip any stationary bandit algorithm with non-stationarity handling capabilities. MASTER [62] is the only known order-optimal black-box method for general bandit and reinforcement learning settings. Importantly, although MASTER is order-optimal, it is not practically applicable [28]. More broadly, the literature emphasizes theory over evidence, as empirical validation of order-optimal methods is scarce: NS-NPBs and NS-PBs are evaluated almost exclusively on synthetic data [61, 30, 28], and NS-CBs lack experiments altogether [15]. We close these gaps with a theoretically grounded, practical black-box framework and comprehensive real-world evaluation in standard benchmarks.
Contributions.
We present (to our knowledge) the first practical prior-free, black-box detection-based framework for general NS bandits. The design is motivated by three pragmatic insights: (i) prior knowledge of non-stationarity is rarely available, (ii) restart-style methods can have lower worst-case complexity than fully adaptive schemes [46], and (iii) a black-box reduction simplifies NS algorithm design to specifying when to restart a stationary learner. Our method is simple—combining a change detector with any stationary bandit algorithm—modular, and easy to implement. Empirically, extensive synthetic and real-world evaluations in standard datasets show consistent gains over both prior-free and prior-based baselines, and (to our knowledge) provide the first comprehensive real-world assessment of order-optimal baselines previously lacking empirical study. Theoretically, under mild assumptions, our regret matches the state-of-the art for PS-LBs, PS-GLBs and PS-CBs and improves the best known bounds for PS-SCBs and PS-KBs; for drifting regimes we identify conditions for good performance and validate them empirically.
@@ 2 The DAL Framework
To identify changes in the environment, DAL uses a general-purpose detector $\mathcal{D}$ for monitoring shifts in the distribution of judiciously chosen reward observation sequences obtained through forced exploration. This distinguishes our approach from methods like MASTER, which rely on detecting violations of stationary regret guarantees. We adopt a detector aligned with [9, 31], grounded in the well-established theory of quickest change detection [60, 65]. Given any arbitrary context, DAL samples rewards from actions within a carefully selected finite subset, and detects changes in the mean reward associated with the context-action pair.
@@
In stationary bandit settings, effective algorithms quickly concentrate on (near-)optimal actions for each context, rarely exploring suboptimal actions. In NS environments, however, this behavior may lead to missed changes on these rarely sampled actions, and thus, forced exploration on these actions is essential. When the action space is large or infinite, exploring all actions becomes infeasible. Therefore, DAL only does extra exploration on a finite covering set, $\mathcal{A}_{\mathrm{e}}=\{a^{(i)}:i\in[N_{\mathrm{e}}]\}\subseteq\mathcal{A}$, in which $a^{(i)}$ denotes the $i$-th action in $\mathcal{A}_{\mathrm{e}}$. $\mathcal{A}_{\mathrm{e}}$ is designed such that the mean reward of at least one context-action pair in $\mathcal{C}\times\mathcal{A}_{\mathrm{e}}$ changes whenever a change occurs. In particular, after the $(k-1)^{\mathrm{th}}$ restart, DAL is forced to play each action in $\mathcal{A}_{\mathrm{e}}$ once for $N_{\mathrm{e}}$ steps, followed by the bandit algorithm for the next $\lceil N_{\mathrm{e}}/\alpha_{k}\rceil-N_{\mathrm{e}}$ steps, repeatedly, until the $k^{\mathrm{th}}$ restart. Here, $\alpha_{k}\in(0,1)$ is the exploration frequency, striking a balance between detection delay and regret from extra exploration.
@@
With a detector $\mathcal{D}$ and forced exploration, DAL augments a (stationary) bandit algorithm $\mathcal{B}$: It resets $\mathcal{B}$ entirely whenever $\mathcal{D}$ detects changes in a reward distribution associated with any context-action pair in $\mathcal{C}\times\mathcal{A}_{\mathrm{e}}$, and runs $\mathcal{B}$ with periodic forced exploration otherwise. A key advantage of DAL is its ability to translate strong stationary performance into robust performance under NS conditions. Therefore, by selecting a well-performing bandit algorithm, the DAL framework inherently achieves effective adaptation to NS environments. In fact, the only requirement for DAL’s input stationary algorithm is to attain optimal stationary regret performance bounds.
@@
Input: bandit $\mathcal{B}$, detector $\mathcal{D}$, covering set $\mathcal{A}_{\mathrm{e}}$, context set $\mathcal{C}$, horizon $T$, frequencies $\{\alpha_{k}\}_{k=1}^{T}$
Initialize: histories $\mathcal{H}_{(c,a)}\leftarrow\emptyset$$\forall(c,a)$$\in\mathcal{C}\times$$\mathcal{A}_{\mathrm{e}}$, detection $\tau$ $\leftarrow$ 0, counter $k$ $\leftarrow$ 11
@@
2:  Observe context $C_{t}$
3:  if $(t-\tau+1\mod\lceil N_{\mathrm{e}}/\alpha_{k}\rceil)$$1=i\in[N_{e}]$ then
4:   Play action $a^{(i)}$ and receive reward $X_{t}$
5:   Add sample $X_{t}$ into history $\mathcal{H}_{(C_{t},a^{(i)})}$
6:   if $\mathcal{D}\left(\mathcal{H}_{(C_{t},a^{(i)})}\right)=\mathrm{detection}$ then
7:    Reset the bandit algorithm $\mathcal{B}$
8:    Clear all $\mathcal{H}_{(c,a)}\kern 5.0pt\forall(c,a)\in\mathcal{C}\times$$\mathcal{A}_{\mathrm{e}}$,
9:    $\tau\leftarrow t$, $k\leftarrow k+1$
10:   end if
11:  else
12:   Run the stationary bandit algorithm $\mathcal{B}$
13:  end if
14:end for
@@ 3 Practical Performance
We evaluate DAL on piecewise-stationarity and drifting non-stationarity across synthetic and real-world benchmarks. Baselines include: MASTER [62], the only other black-box algorithm with order-optimal regret. While MASTER has no formal guarantees for SCBs, empirical evidence [61] supports pairing it with Logistic-UCB-1 [21] for NS-SCBs. We also include two state-of-the-art prior-free, order-optimal methods: ADA-OPKB [30] for NS-LBs/NS-KBs and ADA-ILCTB+ [15] for NS-CBs. ADA-OPKB requires extensive parameter tuning for strong performance (7 hyper-parameters), posing a challenge in a truly prior-free environment. For fair comparison, however, we tune its parameters to optimize its performance in our evaluation. We also tune MASTER’s single hyperparameter ($n$) for best performance. Finally, two prior-based discounted methods are used for comparison: WeightUCB [61] (drifting PBs and PS-SCBs) and WGP-UCB [19] (drifting KBs). Unless noted otherwise, we adopt the hyperparameters in the prior works in our experiments.
Across all settings, DAL uses the Generalized Likelihood Ratio (GLR) test [32] as the detector $\mathcal{D}$. Concretely: In NS-LBs, LinUCB [1] pairs with Gaussian GLR. In NS-GLBs, GLM-UCB [24] pairs with Gaussian GLR. In NS-SCBs, OFUGLB [38] pairs with Bernoulli GLR. In NS-KBs, REDS [52] pairs with Gaussian GLR. In NS-CBs, SquareCB [26] pairs with Bernoulli GLR. The GLR test implementation follows [31], and we provide their structure in the Appendix. For all settings, we set $\alpha_{k}=\sqrt{k||\mathcal{C}|N_{\mathrm{e}}}/(2\sqrt{T}\log^{2}T)$ as per Theorem 4.4. A crucial advantage of DAL is that it is hyperparameter-free, guided entirely by our theoretical principles. The selection of $\mathcal{A}_{\mathrm{e}}$ is discussed more analytically in the Appendix.